OpenStudy (anonymous):

The radius needed to create a sphere with a given volume V can be appromixated by the equation r=0.62(V)^1/3. Find the radius of a sphere with a volume of 125 cubic meters.

OpenStudy (anonymous):

3.1

OpenStudy (anonymous):

\[V=4/3 \pi*r^3\] SO: \[(3V/4\pi)^{1/3}=((3*125)/4\pi)^{1/3}=r=3.1017\]